Time to swing. The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T2 =π^2L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing?
Answer by stanbon(75887)

Time to swing. The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T2 =π^2L.
If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing?
-------
8T^2 = pi^2*10
t^2 = (5/4)pi^2
t = (1/2)*sqrt(5)*pi
t = 3.512 seconds
========================
Cheers,
Stan H.